# D-LIO: 6DoF Direct LiDAR-Inertial Odometry based on Simultaneous Truncated Distance Field Mapping

### v1

@@ V-A VIRAL Dataset
Table I: Experimental results, ATE (m), on VIRAL and New College Dataset across different algorithms. The best result is highlighted in bold and the second best is underlined, excluding LIO-SAM as it performs back-end optimization.
D-LIO A-LOAM M-LOAM FAST-LIO2 KISS-ICP LIO-SAM
eee01 0.092 0.212 0.249 0.166 2.383 0.075
eee02 0.083 0.199 0.168 0.100 1.586 0.069
eee03 0.122 0.148 0.233 0.142 1.055 0.101
nya01 0.080 0.077 0.123 0.127 0.359 0.076
nya02 0.104 0.091 0.191 0.151 - 0.07
nya03 0.065 0.108 0.226 0.130 1.389 0.137
sbs01 0.094 0.103 0.173 0.130 1.353 0.089
sbs02 0.061 0.091 0.147 0.144 1.435 0.083
sbs03 0.098 0.367 0.153 0.126 1.037 0.14
quad-easy 0.093 0.085 0.141 0.100 0.100 0.074
Figure 3 shows the 3D reconstruction generated by D-LIO, along with the trajectory used to obtain it. More comprehensive results, including the estimated trajectories from all runs and their associated visualizations, will be made available in the corresponding GitHub repository [^2] https://anonymous.4open.science/r/D-LIO.
Figure 3: 3D Map reconstruction of maps “eee”,“nya” and “sbs” alongside with each estimated 2D trajectory, where groundtruth is shown in red and the estimated trajectory in blue.
V-B Newer College Dataset
In this dataset, the “Quad-Easy” trajectory has been followed, which involves two loops in the quad area with a nominal walking speed. The platform is equipped with a series of sensors, of which we have only used the LiDAR “Ouster OS-0-128”, which already comes with an integrated IMU, as the inertial information also comes from the IMU integrated into the LiDAR. In this case, the platform is manually carried by a person, who performs a couple of loops over the environment at a constant speed. For these experiments, the map size has been set to $80\times 60\times 35$ meters, and the parameters $\lambda=1.0$, $t_{\text{th}}=1\,\text{m}$, and $q_{\text{th}}=25^{\circ}$.
As shown in Figure 4, the trajectory followed is relatively simple, and the D-LIO successfully estimates a value comparable to those achieved by state-of-the-art methods listed in Table I. The particularity of this environment lies in both its large scale (approximately 80x60x35 meters) and the high point density of the LiDAR sensor, around 100 k points per frame. The objective of testing this dataset is to observe the algorithm’s scalability in environments with high point densities, as existing algorithms using odometry with distance field generation struggle with such scalability.
Figure 4: 3D Map reconstruction of map ”Quad-Easy” from Newer College Dataset alongside with estimated trajectory.
V-C Runtime and memory consumption
The experiments were performed on an HP Victus 16 laptop equipped with 32 GB RAM and a 13th generation Intel Core i7-13700H processor. No GPU was used for the experimentation. The performance of the system was analyzed in terms of memory usage, CPU load, and execution time for the two datasets analyzed, using the same configuration in both cases. Regarding memory consumption, the college dataset exhibits a footprint ranging from 30% to 90%, while in the VIRAL dataset, memory usage is generally lower, between 10% and 40%. In both cases, the grid update step is the primary contributor to memory usage.
Overall memory consumption is primarily determined by the size of the selected map, since a fixed portion of memory is reserved for the grid. Consequently, the experiments on the college dataset consumed around 19 GB of memory, while those on VIRAL, with a smaller map, required approximately 10 GB.
Execution times are summarized in Table II, reporting the average times measured over several trajectories without downsampling, thus reflecting the most demanding scenario. The map update step consistently emerges as the most computationally expensive stage, often becoming the bottleneck of the algorithm. This update time strongly depends on the density of the point cloud to be integrated. For instance, in the VIRAL dataset, the lower density of points from both the horizontal and vertical OS-1 (approx. 20 k points per scan) leads to shorter map update times ($\approx 0.30$ s), whereas in the college dataset, where the LiDAR produces around 100 k points/scan, the update step is significantly slower ($\approx 1.2$ s). However, the map update time has little impact on the average total time (see Table II) because our method performs the update operation at a much lower frequency than the sensor acquisition (10 Hz), while the optimization step, which is executed at every scan, remains lightweight (typically below 0.05 s). This design allows the overall average computation time per scan to remain below the LiDAR rate in most cases, except for the College dataset, where the high point density and frequent updates during longer motion intervals result in a higher total runtime.
@@ V-C Runtime and memory consumption
It is also worth noting that state-of-the-art distance field mapping approaches, such as VDB-GPDF [14], which focuses solely on reconstruction, reported similar execution times to ours, but our approach performs both odometry and reconstruction without any downsampling. Although the map is updated less frequently, once every meter, it integrates the full-resolution point cloud, offering a balanced trade-off between computational efficiency and mapping accuracy.
@@
Table II: Runtime analysis across different datasets, showing mean ± standard deviation (in seconds) for each processing phase.
Dataset Total (s) Optimize (s) Update (s)
college 0.583 ± 0.433 0.430 ± 0.187 1.216 ± 0.060
eee 0.052 ± 0.051 0.040 ± 0.015 0.301 ± 0.036
nya 0.083 ± 0.071 0.065 ± 0.024 0.367 ± 0.022
sbs 0.051 ± 0.048 0.044 ± 0.016 0.301 ± 0.014
@@ VI CONCLUSIONS AND FUTURE WORK
D-LIO, a fully direct LiDAR–inertial odometry pipeline based on Fast-TDF simultaneous mapping, has been shown to match or outperform feature-driven state-of-the-art methods while producing a continuously available distance field for downstream tasks. Unlike conventional TSDF/ESDF frameworks whose update cost grows with map volume, our Fast-TDF binary-mask scheme ties each integration step solely to the number of incoming points. By encoding obstacle proximity directly in the truncated distance field, we avoid feature extraction, yielding robust convergence even in feature-sparse or highly reflective environments. The resulting TDF is immediately usable for collision avoidance or planning without any post-processing.
Future work will focus on two main extensions. First, addressing the limitation of using a fixed-size grid, as there is currently no strategy to manage the map when the system approaches its spatial limits. Introducing dynamic resizing or shifting methods could improve scalability in larger environments. Second, extending the Fast-TDF kernel to support signed distance estimation, enabling the system to distinguish whether a point lies in front of or behind an obstacle, potentially enhancing registration accuracy.

### v2

@@ V-A Trajectory evaluation
Table I: Experimental results on VIRAL and New College Dataset: ATE (m) for all methods (m). The best result is highlighted in bold and the second best is underlined, excluding LIO-SAM and PIN-SLAM as they perform back-end optimization. D-LIO, FAST-LIO2, and LIO-SAM are LiDAR–inertial methods, while the others rely solely on LiDAR. Metrics were taken from [40], [42] and [41].
D-LIO A-LOAM M-LOAM KISS-ICP FAST-LIO2 LIO-SAM PIN-SLAM
eee01 0.074 0.212 0.249 2.383 0.166 0.075 -
eee02 0.073 0.199 0.168 1.586 0.100 0.069 2.056
eee03 0.121 0.148 0.233 1.055 0.142 0.101 0.615
nya01 0.080 0.077 0.123 0.359 0.127 0.076 0.086
nya02 0.104 0.091 0.191 - 0.151 0.07 0.092
nya03 0.065 0.108 0.226 1.389 0.130 0.137 0.455
sbs01 0.085 0.103 0.173 1.353 0.130 0.089 0.205
sbs02 0.067 0.091 0.147 1.435 0.144 0.083 0.528
sbs03 0.098 0.367 0.153 1.037 0.126 0.14 0.708
q-easy 0.093 0.085 0.141 0.100 0.100 0.074 0.09
Figure 3: 3D map reconstructions (from left to right) of “eee”, “nya”, “sbs”, and Newer College (Quad-Easy) datasets.
V-B Scalability
After validating the accuracy of our algorithm in more dynamic trajectories, we next assess scalability on longer trajectories and larger maps using the VBR Dataset (See Table II). On the “diag” sequence, we allocated 250k high-resolution blocks within a $\times150$$\times$50 m map. Despite the moderate map size, the trajectory is long (1.4 km); however, the points are compactly distributed, so this allocation sufficed to represent the space without reallocations, as cells are created only around observed points, avoiding empty space and improving memory efficiency. Results are competitive with the state of the art, second only to PIN-SLAM, which benefits from global loop closures on long, loopy paths. In the “pincio” sequence, the platform covers 2.5km over a much wider area. Since the full map cannot be kept in memory without reallocations, we limited the allocation to 50k high-resolution blocks. The system preserves detail only near the current pose, dynamically evicting distant blocks that no longer affect the current scan, thus managing memory dynamically. As this run spans a larger area and lacks loop closures, the absolute error increases; nevertheless, an ATE of 7 m over 2.5 km corresponds to only 0.28% of the total trajectory length.
Table II: Experimental results on VBR Dataset: ATE (m) and RPE (%) in parenthesis for all methods. The best result is highlighted in bold and the second best is underlined.
D-LIO KISS-ICP PIN-SLAM
diag 0.560 (0.72) 1.397 (1.790) 0.362 (0.468)
pincio 7.39 (3.45) 0.784 (0.485) 0.647 (0.453)
V-C Mapping
To quantitatively evaluate the mapping performance of our algorithm, we adopt standard metrics widely used for 3D reconstruction benchmarking, following the evaluation methodology proposed in [43]. The evaluation includes four key indicators: Chamfer-L1 distance (C-L1), reconstruction accuracy, completeness, and F-Score. Chamfer-L1 offers an aggregate measure integrating both reconstruction accuracy, representing the fidelity to the ground truth, and completeness, which reflects how thoroughly the environment has been captured. For accuracy, completeness, and Chamfer-L1 metrics, smaller values represent improved map quality. Conversely, higher F-Score values reflect a better balance between precision and recall. The detailed results of these metrics are summarized in Table III
Table III: Mapping metrics on Newer College Quad-Easy sequence. Best and second best scores are marked in bold and underlined respectively
Pose Acc. Comp. C-L1 F-Score
VDB-Fusion KissICP 14.03 25.46 19.75 69.50
SHINE 14.87 20.02 17.45 68.85
NKSR 15.67 36.87 26.67 58.57
PUMA Odometry 15.30 71.91 43.60 57.27
SLAMesh 19.21 48.83 34.02 45.24
NeRF-LOAM 12.89 22.21 17.55 74.37
S²KAN-SLAM 13.32 18.80 16.06 72.03
ImMesh 15.05 19.80 17.42 66.87
PIN-SLAM 11.55 15.25 13.40 82.08
KN-LIO 8.18 11.65 9.92 91.01
D-LIO 12.14 10.63 11.39 83.83
Overall, D-LIO remains at the state of the art: it achieves the best completeness (10.63), second-best Chamfer-L1 (11.39) and F-Score (83.83), and competitive accuracy (12.14). It approaches KN-LIO’s top performance and surpasses PIN-SLAM on Chamfer-L1, completeness, and F-Score, while classical TSDF/SDF and several odometry-driven baselines lag behind.
V-D Runtime and memory consumption
The performance of the system was analyzed in terms of memory usage, CPU load, and execution time for all datasets. Overall memory consumption is governed by the number of high-resolution blocks (each a $1\,\text{m}^{3}$ region discretized at 0.05 m) rather than the coarse map, whose footprint is negligible by comparison (see Sec. V-A). In VIRAL (largest trajectory), we used 180k blocks ($31\%$ of RAM). In Newer College, due to the higher point concentration, only 45k blocks ($22\%$ of RAM) were required. On diag, despite a 1.4 km path within one building, most points fell into already-activated regions; with an intentionally oversized budget of 250k blocks ($47\%$ of RAM) we needed no reallocations. Finally, on pincio, points were widely dispersed with little coarse-block reuse, so we ran a smaller allocation of 45k blocks ($22\%$ of RAM) with dynamic reallocation as the platform advanced.
Execution times are summarized in Table IV. The map update step normally emerges as the most computationally expensive stage, as its time strongly depends on the point cloud density. For instance, in the VIRAL dataset, the lower density from OS-1 sensors ($\approx$ 20 k points/scan) leads to shorter update times, whereas in the college and VBR dataset, the OS-0 sensor produces $\approx$ 100-260 k points/scan making updates slower. However this cost also depends on environment topology: in diag, compact distributions and frequent revisits mean many already-seen regions are not updated, so the update time can match, or even be lower than, the optimization.
Additionally, updates are triggered at a much lower frequency than optimization (which runs at the sensor rate, 10 Hz), so the weighted average tends to be driven by the optimization step, also influenced by the point density. Note that all results were obtained under a worst-case computational load, no downsampling; in practice, applying voxel filtering can significantly speed up the system with negligible impact on accuracy.
@@ V-D Runtime and memory consumption
It is also worth noting that state-of-the-art distance field mapping approaches on CPU, such as VDB-GPDF [13], which focuses solely on reconstruction, reported similar execution times to ours, but our approach performs both odometry and reconstruction without any downsampling. Although the map is updated less frequently, once every meter, it integrates the full-resolution point cloud, offering a balanced trade-off between computational efficiency and mapping accuracy.
@@
Table IV: Runtime analysis across different datasets with no downsample, showing mean ± standard deviation (in seconds) for each processing phase. Also reported, per trajectory, the average count of truly measured LiDAR returns.
Points Total (s) Optimize (s) Update (s)
eee 20k 0.048 ± 0.055 0.046 ± 0.013 0.155 ± 0.009
nya 20k 0.049 ± 0.034 0.042 ± 0.014 0.146 ± 0.016
sbs 20k 0.051 ± 0.048 0.044 ± 0.016 0.142 ± 0.034
q-easy 100k 0.479 ± 0.196 0.442 ± 0.158 0.470 ± 0.095
diag 190k 0.603 ± 0.240 0.565 ± 0.194 0.315 ± 0.085
pincio 140k 0.576 ± 0.256 0.505 ± 0.182 0.563 ± 0.146
@@ VI CONCLUSIONS AND FUTURE WORK
D-LIO, a fully direct LiDAR–inertial odometry pipeline based on Fast-TDF simultaneous mapping, has been shown to match or outperform feature-driven state-of-the-art methods while producing a continuously available distance field for downstream tasks. Unlike conventional Neural TSDF/ESDF frameworks whose update cost grows with map volume, our Fast-TDF binary-mask scheme ties each integration step solely to the number of incoming points. By encoding obstacle proximity directly in the truncated distance field, we avoid feature extraction, yielding robust convergence even in feature-sparse or highly reflective environments.
Future work will focus on extending the Fast-TDF kernel to support signed distance estimation, enabling the system to distinguish whether a point lies in front of or behind an obstacle, potentially enhancing registration accuracy.